OpenStudy (anonymous):

solve for x: [x-2] + 10 = 12 x= 0 and x = 4 x= -4 and x= 0 x= -20 and x= 4 no solution

OpenStudy (anonymous):

@divu.mkr

OpenStudy (austinl):

[x-2]+10=12 or |x-2|+10=12

OpenStudy (anonymous):

A if its a mod.. (:

OpenStudy (austinl):

Based on the answers it would be my second choice above. So we have this, \(|x-2|+10=12\) Subtract 10 from both sides, \(|x-2|=2\) Which is the same as, \(-(x-2)=2\\ \&\\ x-2=2\) Solve for x in each of those and it is your answer!

OpenStudy (austinl):

@pic1e have you arrived at an answer yet?

OpenStudy (anonymous):

no

OpenStudy (austinl):

Solve, -(x-2)=2 and x-2=2 for x. You will have two answers. What do you get when you do this?

OpenStudy (anonymous):

@divu.mkr help me pleaseee

OpenStudy (anonymous):

u want a solution?

OpenStudy (anonymous):

yes @divu.mkr

OpenStudy (anonymous):

then mod is defined as x-2=2 if x>=2 and -(x-2) =2 that will give you 0,4

OpenStudy (anonymous):

omg thankyou !!!!!!! can u help me with some more ? @divu.mkr

OpenStudy (anonymous):

yep..